Every year , the population of a certain country increases by 2%. If the population is now 20 million people, what will be the population after 8 years?


1
Expert's answer
2021-09-07T18:20:48-0400

Given, the population P=20,000,000,the rate of increaseing R=2% and the number of years n=8yr.


formula,

Population after n years = "P(1+\frac{R}{100})^n\\
=20000000(1+\frac{2}{100})^8\\
=23433187.62"